OpenStudy (anonymous):

1. Determine the zeros of f(x) = x3 – 3x2 – 16x + 48? Somebody explain it please?:/

OpenStudy (anonymous):

OH! I already know that there are 2or0 positive, 1 negative and 2or0 complex. But I dont know what to do next?

OpenStudy (anonymous):

f(x) = x3 – 3x2 – 16x + 48 =x^2(x-3)-16(x-3) =(x-3)(x^2-16) = (x-3)(x-4)(x+4) for f(x) =0 f(x) =(x-3)(x-4)(x+4) =0 x =3 , 4 , -4

OpenStudy (lgbasallote):

are you familiar with rational roots theorem?

OpenStudy (anonymous):

that the answer for all

OpenStudy (anonymous):

...@abayomi12 that made no sense to me. I would like an explanation and @lgbasallote no I am not..

OpenStudy (lgbasallote):

hmm...im trying to think of a way to solve this without using rational roots theorem. are you sure it was never taught to you before?

OpenStudy (anonymous):

I'm doing virtual school. The online lesson isn't very clear about it. So if you have any other way to find the zero's that would be great!

OpenStudy (lgbasallote):

well let me tell you about the general concept first.. the steps to finding the zeros of CUBIC functions are these: Step 1: Find ONE factor of the expression Step 2: Divide the expression by that factor (the quotient will be a quadratic equation) Step 3: Factor or use the Quadratic Formula on that quadratic equation to get hte remaining roots the rational roots theorem is used to find one of the factors for step 1

OpenStudy (lgbasallote):

now as for an alternative method to rational roots theorem...i would have to ask for a consult from one of my colleagues.. @waterineyes would you be kind enough to grace us with your wisdom?

OpenStudy (amistre64):

abayomi has the right idea in this case, try to factor the poly

OpenStudy (amistre64):

try and error different combonations that have common factors to pull out (x3 – 3x2) – 16x + 48? x^2(x-3) -16x+48 it would be great of the (x-3) part is a factor of whats left on the right n(x - 3) xn - 3n = -16x+48 +16x +16x -------------------- (n+16)x - 3n = 48 +3n +3n --------------------- (n+16)x = 48 + 3n = 0 48 + 3n = 0 3n = -48 n = -16 if (-16+16)x = 0 , then we are good to go with this factoring ... most times this is done in the head tho